OpenStudy (anonymous):

What is the solution set of the given system? 2y = x + 3 5y = x - 7

OpenStudy (anonymous):

Answer choices; {(-1/3, -4/3)} {(1/3, 4/3)} {(-29/3, -10/3)} {(29/3, 10/3)} Someone Help Please

OpenStudy (anonymous):

its the second last one

OpenStudy (anonymous):

{(-29/3, -10/3)}

OpenStudy (anonymous):

third one. 2y=x+3 x = 2y - 3 Substituting this in second equation: 5y = 2y - 3 -7 3y = -10 y = -10/3. So\[x = (2\times\frac{-10 }{ 3 }) -3 = = \frac{ -20 }{ 3 } - 3 = \frac{ -29 }{ 3 }\]

OpenStudy (anonymous):

So its third one.